Lily Company has the following data for the weekly payroll ending January 31.

Hours

Employee

M

T

W

T

F

S

Hourly Rate

Federal Income Tax Withholding

Health Insurance

L. Helton 9 8 8 8 10 2 $12 $35 $11
R. Kenseth 7 9 8 7 7 4 14 37 26
D. Tavaras 9 11 9 9 9 0 15 58 25


Employees are paid 1.5 times the regular hourly rate for all hours worked in excess of 40 hours per week. FICA taxes are 7.65% on the first $127,200 of gross earnings. Lily Company is subject to 5.4% state unemployment taxes and 0.8% federal unemployment taxes on the first $7,000 of gross earnings.

a. Prepare the payroll register for the weekly payroll. (Round answers to 2 decimal places, e.g. 15.25.)

LILY COMPANY
Payroll Register
For the Week Ending January 31

Earnings

Deductions

Employee

Total Hours

Regular

Overtime

Gross Pay

FICA Taxes

Federal Income Taxes

Health Insurance

Total

Net Pay

L. Helton

$

$

$

$

$

$

$

$

R. Kenseth
D. Tavaras

Totals

$

$

$

$

$

$

$

$

b. Prepare the journal entries to record the payroll and Lily’s payroll tax expense. (Credit account titles are automatically indented when amount is entered. Do not indent manually. Round answers to 2 decimal places, e.g. 15.25.)
Answer #1
Lily Company
Payroll Register
For the Week ending Jan.31
Earnings Deductions Paid
Employee Total Hours Regular
(a)
(40 HourX Hourly rate)
Over-Time
(b)
Gross
(c=a+b)
FICA
(d=c*7.65%)
Federal I Tax Withholding
(e )
Health Insurance (f) Total
(g=d+e+f)
Net Pay
(c-g)
M Pinson 45 $480.00 $90.00 $570.00 $43.61 $35.00 $11.00 $89.61 $480.40
E. Tavaras 42 $560.00 $36.00 $596.00 $45.59 $37.00 $26.00 $108.59 $487.41
K. Helton 47 $600.00 $126.00 $726.00 $55.54 $58.00 $25.00 $138.54 $587.46
Total $1,640.00 $252.00 $1,892.00 $144.74 $130.00 $62.00 $336.74 $1,555.26
Journal Entry
Date Account Tittle Debit Credit
31-Jan Office Salaries $1,892.00
FICA $144.74
Fedearl Tax Withholding $130.00
Health Insurance $62.00
Office Salaries Payable A/c $1,555.26
31-Jan Payroll Tax Exp. $117.30
State Unemployment Tax
(1892*5.4%)
$102.17
Federal Unemployment Tax
(1892*0.8%)
$15.14